OpenStudy (anonymous):

Solve the system of equations 7x+3y=30 -2x+3y=3

OpenStudy (anonymous):

ok lets use the elimination method since we have 3y and 3y in the first and last one already, we can subtract directly. 7x -(-2x) 3y - 3y = 30 -3 9x = 27 x = 27/9 x = 3

OpenStudy (anonymous):